# Scalable Hybrid quantum Monte Carlo simulation of U(1) gauge field coupled to fermions on GPU

### v1

@@ Abstract
We develop a GPU-accelerated hybrid quantum Monte Carlo (QMC) algorithm to solve the fundamental yet difficult problem of $U(1)$ gauge field coupled to fermions, which gives rise to a $U(1)$ Dirac spin liquid state under the description of (2+1)d quantum electrodynamics QED3. The algorithm renders a good acceptance rate and, more importantly, nearly linear space-time volume scaling in computational complexity $O(N_{\tau}V_{s})$, where $N_{\tau}$ is the imaginary time dimension and $V_{s}$ is spatial volume, which is much more efficient than determinant QMC with scaling behavior of $O(N_{\tau}V_{s}^{3})$. Such acceleration is achieved via a collection of technical improvements, including (i) the design of the efficient problem-specific preconditioner, (ii) customized CUDA kernel for matrix-vector multiplication, and (iii) CUDA Graph implementation on the GPU. These advances allow us to simulate the $U(1)$ Dirac spin liquid state with unprecedentedly large system sizes, which is up to $N_{\tau}\times L\times L=660\times 66\times 66$, and reveal its novel properties. With these technical improvements, we see the asymptotic convergence in the scaling dimensions of various fermion bilinear operators and the conserved current operator when approaching the thermodynamic limit. The scaling dimensions find good agreement with field-theoretical expectation, which provides supporting evidence for the conformal nature of the $U(1)$ Dirac spin liquid state in the QED3. Our technical advancements open an avenue to study the Dirac spin liquid state and its transition towards symmetry-breaking phases at larger system sizes and with less computational burden.
I Introduction
I.1 General Background
The emergence of Dirac spin liquid (DSL), i.e. $U(1)$ gauge field coupled to gapless relativistic spinons at (2+1)d [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], or, in high-energy physics terms, a deconfined phase in 3D quantum electrodynamics (QED3) coupled to $N_{f}$ flavor massless fermions [13, 14, 15, 16, 17, 18, 19, 20], has hold the attention of condensed matter and high-energy physicists for decades. QED3 is expected to realize a conformal field theory (CFT) for large $N_{f}$ [19, 17], which contains many interesting scaling operators including fermion bilinears operators in the form of $SU(N)$ flavor currents and various fermion bilinear terms, as well as monopole operators, which generate $U(1)$ magnetic fluxes and transform non-trivially on lattice.
Over the years, the QED3 CFT becomes the prototypical problem for testing perturbative and non-perturbative methods to study strongly coupled conformal field theory in 2+1 dimensions in the high energy community [13, 14, 15, 16, 19, 18, 17, 20]. In condensed matter physics, QED3 state has been argued to arise as a ground state of certain frustrated quantum spin models with Heisenberg-type Hamiltonian, with no magnetic order (no spontaneous symmetry breaking) and yet the emergent spinon exhibiting Dirac dispersion and their interactions mediated by the fluctuating $U(1)$ gauge field. It is in this context such a gapless and critical state is known as $U(1)$ DSL [21, 1, 2, 3, 4, 5, 22, 23, 24, 25, 26, 27, 28, 29, 6, 7, 8, 9, 30, 31, 32, 10, 11, 12, 33, 34].
From a field-theoretical perspective, the basic properties of QED3 are understood, e.g. one knows the primary fields and their roughly estimated scaling dimensions [18, 17, 20]. However, concrete lattice model calculations, with the precise properties of multi-point correlation functions and their evolution across phase transition to other symmetry-breaking phases, tuned by quantum and thermal knobs such as the fluctuation strength, magnetic field, temperature, etc, are yet unknown. This, to the authors viewpoint, is mainly due to the lack of controlled computational scheme that can solve this strongly interacting theory in an unbiased manner and extrapolate, for example, the correlation functions of various fermion bilinears, from finite space-time volume to the thermodynamic limit. It is such difficulty that motivates us to develop the scalable hybrid quantum Monte Carlo on GPU to solve this problem in this work.
At the condensed matter experimental front, the signature of (proximate) $U(1)$ DSL has been actively reported in frustrated quantum magnets in recent years, in particular triangular lattice antiferromagnets YbZn2GaO5 [35, 36] and the $A$-YbSe2 delafossites [37, 38] and kagomé antiferromagnets [39, 40, 41, 42, 43, 44, 45]. For example, in the recently synthesized kagomé material YCu3(OD)6Br2[Br1-x(OD)x], thermodynamic measurements have revealed clear $T^{2}$ dependence in specific heat, suggesting gapless degrees of freedom with linear dispersion in this electronically gapped Mott insulator [39]. Subsequently, inelastic neutron scattering reveals spin spectral signature of the Dirac cone continuum for a nonmagnetic ordered state at low temperature [40, 41], and magnetic torqe measurements found unconventional quantum oscillations near the 1/9 magnetization plateau in this Mott insulator [45], all these results are pointing towards a DSL state in the material.
In light of these progresses, it is timely to develop a quantum many-body computation scheme that can access the QED3 $U(1)$ DSL state with ease and control, such that one can address and understand the theoretical questions and make ossible connections to experimental observations.
I.2 Computational Challenges
As mentioned above, theoretical approaches cannot arrive at controlled results unless $N_{f}$ is large, but the physical situation is at $N_{f}=2$ for spin-1/2 fermions. Considering the fermion doubling on the lattice, there will be $2N_{f}$ Dirac fermions and the lattice symmetry is enlarged to an emergent $SU(2N_{f})=SU(4)$ symmetry. In terms of concrete lattice model calculations, only very recently, unbiased numerical results of QED3 on finite size lattices from determinant quantum Monte Carlo (DQMC) simulations start to appear [6, 7, 9], but the simulation can not access sufficiently large system sizes and low temperature as it suffers heavily from unfriendly computational complexity of fermion determinant and slow Monte Carlo dynamics due to local update in the large configurational space spanned by the continuous $U(1)$ variables in (2+1)d space-time lattice.
To be more specific, even with specially designed fast updates (it is still a local update in the gauge field configuration) [6], the computational complexity of DQMC remains challenging at $O(N_{\tau}V_{s}^{3})$, with $N_{\tau}$ the imaginary time dimension and $V_{s}=L^{2}$ the number of spatial sites ($L$ is the linear dimension). The primary source of this scaling stems from the explicit determinant evaluation in the partition function. Calculating the acceptance ratio in DQMC depends on the equal-time Green’s function, and updating the Green’s function after sweeping through the space-time lattice renders a complexity of either $O(N_{\tau}V_{s}^{3})$ (with fast updates) or $O(N_{\tau}V_{s}^{4})$ (without fast updates).
Moreover, the $U(1)$ DSL phase is a critical phase characterized by gapless dynamics of both the gauge field and the fermion. In such phase, the simulation usually experiences critical slowing down, i.e. the Monte Carlo autocorrelation time $\tau_{\text{int}}$ of physical observables grows with system size as $\tau_{\text{int}}\sim L^{z_{\text{MC}}}$, where $z_{\text{MC}}$ is the algorithmic dynamical exponent reflecting the efficiency of the Monte Carlo update scheme. DQMC with Metropolis update for critical phases often renders a larger $z_{\text{MC}}$ than its classical counterpart ($z_{\text{MC}}=2.17$ for 2D classical Ising model at its critical point) [46, 47]. This inefficiency becomes severe for the continuous gauge field with highly anisotropic terms in the Lagrangian in Eqs. (1), (2) and (3), as the ratio between temporal and spatial terms is of order $\frac{1}{\Delta\tau^{2}}$, with $\Delta\tau$ being the Trotter step size. The dynamics of local update is thus dominated by the temporal term in Lagrangian. More specifically, local updates of the form $\phi^{\prime}_{ij,\tau}=\phi_{ij,\tau}+\Delta\phi$ suffer from low acceptance rates when the gauge coupling $J$ is small—typically in the $U(1)$ DSL phase. Since the energy difference of such updates is inversely proportional to $J$, a smaller $J$ requires a smaller $\Delta\phi$ update to maintain a reasonable acceptance ratio. And a smaller $\Delta\phi$ in turn results in larger autocorrelation time in the sampled configurations.
Therefore the DQMC seems difficult to handle the problem, whether there are other Monte Carlo methods that could come to help? The answer is yes and it is the hybrid quantum Monte Carlo (HQMC) method [48, 49, 50, 51, 52, 53]. At least on paper, it enjoys two clear advantages over DQMC in studying this problem: (i) instead of the local update in DQMC, the Hamilton’s equation of motion in HQMC updates the gauge field configurations globally, while preserving a good acceptance rate; (ii) with the help of psudo-fermion transformation, the evaluation of the fermionic determinant, which has $O(N_{\tau}V_{s}^{3})$ scaling, is replaced by solving a linear system with iterative methods, which can have as low a computation complexity as $O(N_{\tau}V_{s})$, and is also amenable to parallel computing like GPU acceleration. So, in principle, HQMC is a much more efficient method than DQMC.
However, actual application of the HQMC in condensed matter models faces many challenges, which are often problem specific. Previous attempts of employing HQMC on the Hubbard model did not show significant advantages  [48, 54]. In particular, in a recent attempt [53], while the ergodicity issue is successfully solved by a compactification method, the overall time complexity of the HQMC algorithm is not much better than conventional DQMC, because of the diverging iteration number complexity of the linear equation solver. Therefore, doubts are cast upon the applicability of HQMC on strongly correlated systems in condensed matter physics.
But the more recent developments are very encouraging. The applications of HQMC to the spin-fermion models for the investigation of the non-Fermi-liquid and quantum critical metallic state [55, 56], apparently enjoy successes, both in the physical discoveries and in computation technologies. In the spin-fermion models, the bosonic fields originate from the physical spin degrees of freedom which provide the critical fluctuations [57, 58, 59], and are different from the auxiliary bosonic field in Hubbard model which comes from the Hubbard-Stratonovich transformation while decoupling the density-density interaction. It is found that by applying the HQMC on spin-fermion models with various optimization techniques e.g. preconditioned conjugate gradient (pCG) method, matrix-free multiplication and GPU parallel computing, the authors of Ref. [55, 56] significantly scale up the simulation size, which is up to $N_{\tau}\times L\times L=200\times 80\times 80$. The overall computational complexity is also reduced to be linear with respect to time $N_{\tau}$ and spatial volume $V_{s}$, i.e. $O(N_{\tau}V_{s})$.
I.3 Our GPU-accelerated HQMC solution
Inspired by such development. In this paper, we apply HQMC to simulate a $U(1)$ gauge field model coupled to $N_{f}=2$ fermions at (2+1)d space-time on GPU, and manage to scale the simulation size up to $N_{\tau}\times L\times L=660\times 66\times 66$, which, to the best of our knowledge, is a huge improvement in simulation size compared to all previous DQMC studies of this model [6, 7, 9], which is typically up to $N_{\tau}\times L\times L=200\times 20\times 20$, as well as three times larger than the HQMC study of a different spin-fermion model in Ref. [55, 56]. Such improvements are achieved by the implementation of the following several GPU related optimization techniques tailored to this problem, including
  1. 1.
    the problem-specific preconditioner computation,
  2. 2.
@@ I.3 Our GPU-accelerated HQMC solution
    customized CUDA kernel for matrix-vector multiplication that exploits the specific sparsity pattern, and
  3. 3.
    CUDA graph acceleration that minimizes kernel launch overhead.
With these advances, in this fundamental model, computational complexity is now successfully reduced from DQMC’s $O(N_{\tau}V_{s}^{3})$ to a linear complexity $O(N_{\tau}V_{s})$ with respect to imaginary time $N_{\tau}$ and spatial volume $V_{s}$.
Using such GPU-accelerated HQMC, we are able to investigate the existence of $U(1)$ DSL state in QED3 model with ease and control. We study the correlation function of different fermion bilinear operators in DSL state such as spin correlation functions and bond correlations, and find they have the same scaling dimension at the thermodynamic limit, as they belong to the fermion bilinear terms (adjoints) of the emergent $SU(4)$ group and share the same scaling dimension of $\Delta_{\rm adj}\in(1.4,1.7)$ [3, 18]. We reveal such a property for the case of $N_{f}=2$ ($N=4$ due to fermion doubling in DSL phase) QED3, by first benchmarking with DQMC at small system sizes and then performing large-scale simulation up to to $660\times 66\times 66$. We also observe that the flux-flux correlations is dictated by conserved current scaling dimension $\Delta_{J}=2$ of the $SU(4)$ group, such that the long-time correlations of the local flux approach $1/\tau^{4}$ asymptotically. We believe this work opens an avenue to explore the generic phase diagram of the model and the novel phase transitions between $U(1)$ DSL and other phases, in a computationally efficient and scalable manner.
II Models and Methods
II.1 Models
@@
Figure 1: Lattice model of QED3. Fermions live on the sites of the lattice represented by blue dots, while gauge fields live on the bonds of the lattice represented by yellow dots. Fermion hops between nearest-neighbor sites with a phase $e^{i\phi_{ij}}$. The plot shows two adjacent square lattice layers, which are connected with temporal bonds represented as black dashed line. Gauge fields on these temporal bonds are set to be zero. Within a square lattice layer, the four bond colors represent the four families of bonds, which will be used in the checkerboard decomposition, as shown in Eq. (19); within each family, the fermion hopping terms represented by the bonds all mutually commute. In the bottom right cube, the circles on the facets denote the positive directions of the magnetic flux going out of the cube.
@@
The models we study are quantum electrodynamics (QED3) defined on a cubic space-time lattice (square spatial lattice), coupled with two-flavor fermions, as shown in Fig. 1 [6, 7]. Here, we start with the action of these lattice QED3 models: $S=S_{F}+S_{B}=\int_{0}^{\beta}d\tau\left(L_{F}+L_{B}\right)$, where $\beta$ is the inverse temperature and $\tau$ is the imaginary time. The first term
@@
In this paper, we consider two $U(1)$ gauge field Lagrangians, namely compact $U(1)$ gauge field
$L_{B}^{\textmd{c}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left[1-\cos\left(\phi_{ij,\tau+1}-\phi_{ij,\tau}\right)\right]+K\sum_{\square}\cos(\operatorname{curl}\phi),$ (2)
and non-compact $U(1)$ gauge field

### v3

@@ Abstract
The problem of a $U(1)$ gauge field coupled to fermions in (2+1) dimensions is of fundamental importance, as it is believed to give rise to $U(1)$ Dirac spin liquid (DSL) state in quantum magnets and the strongly coupled conformal field theory in quantum electrodynamics. However, numerical progress has long been hampered by the steep computational cost of traditional determinant quantum Monte Carlo (QMC). Here, we develop a GPU-accelerated hybrid QMC algorithm assisted with several novel technical improvements, which significantly reduces the complexity to be linear with respect to space-time volume, and scales the simulation up to an unprecedented size. At this scale, we observe asymptotic convergence of fermion bilinear correlator and conserved current correlator, which was unclear previously, and find their scaling dimensions in good agreement with field theory, which supports the conformal nature of the Dirac spin liquid. Our advances establish a scalable framework to study DSL physics and its transitions at larger scales.
Introduction
The emergence of Dirac spin liquid (DSL), i.e. an emergent $U(1)$ gauge field coupled to gapless relativistic spinons in (2+1)d [Hermele2004, Assaad2005, Hermele2005, ranProjected2007, ranSpontaneous2009, XYXuU12018, wangDynamics2019, LuoTwistedDSL22, WietekQED3, SeifertSpinPeierls24, janssenConfinement2020], or, in high-energy physics terms, a deconfined phase in 3D quantum electrodynamics (QED3) with $N_{f}$ flavors of massless fermions [Fiebig1990, Herbut2003, Fiore2005, Armour2011, chester2016towards, di2017scaling, karthikNumerical2019, Albayrak2022], has attracted continuous attention from condensed matter and high-energy physicists over the past decades. For sufficiently large $N_{f}$, QED3 is expected to flow to a conformal field theory (CFT) in the infrared [karthikNumerical2019, chester2016towards], containing a variety of scaling operators including fermion bilinear currents, mass operators, and monopole operators. Such a critical state has been argued to emerge in frustrated spin systems, where no magnetic order appears and the spinons acquire Dirac dispersions, and their interactions are mediated by fluctuating $U(1)$ gauge fields [Hastings2003, Hermele2004, Assaad2005, Hermele2005, ranProjected2007, ranSpontaneous2009, Iqbal2011, Imada2014, White2015, Hu2015, Iqbal2016, He2017, Gong2017, Saadamand2017, XYXuU12018, wangDynamics2019, song19, He2019, Ferrari2019, nomuraDirac2021, LuoTwistedDSL22, WietekQED3, SeifertSpinPeierls24, Sherman2023, Drescher2023, janssenConfinement2020]. In this condensed matter context, the gapless and critical ground state is referred to as the $U(1)$ DSL.
From the field-theoretical viewpoint, the basic structure of QED3 CFT is understood—one knows the set of primary operators and rough estimates of their scaling dimensions [di2017scaling, chester2016towards, Albayrak2022]. Such quantitative information about the scaling dimensions of fermion bilinear operators is important, since the scaling dimensions govern the asymptotic decay of correlation functions and diagnose the conformal nature of the DSL. In particular, the spin, bond, and dimer correlations correspond to different fermion bilinears that transform in the adjoint representation of the emergent $SU(4)$ symmetry for $N_{f}=2$, and are expected to share the same scaling exponent [rantner2002spin, Hermele2005]. Establishing this universality and comparing with analytical approaches such as large-$N_{f}$ expansion or $\epsilon$-expansion [di2017scaling, chester2016towards] is crucial to confirm the conformal fixed point.
On the condensed matter experimental side, signatures of proximate DSLs have been reported in frustrated magnets, including triangular lattice antiferromagnets YbZn2GaO5 [BagEvidence24, WuSpinDynamics25] and A-YbSe2 delafossites [ScheieKYbSe2, ScheieNaYbSe2], as well as kagomé antiferromagnets [zengPossible2022, zengSpectral2024, hanSpin2024, suetsuguEmergent2024, suetsuguGapless2024, jeonOne2024, zhengUnconventional2025]. For example, the recently synthesized kagomé compound YCu3(OD)6Br2[Br1-x(OD)x] shows a clear $T^{2}$ contribution in specific heat [zengPossible2022], inelastic neutron scattering reveals a continuum consistent with Dirac cones [zengSpectral2024, hanSpin2024], and magnetic torque detects unconventional oscillations near the 1/9 plateau [zhengUnconventional2025]. These results point toward DSL behavior, but a quantitative comparison requires reliable knowledge of the scaling of correlation functions.
On the computational side, progress has been hindered by the absence of efficient and unbiased methods. Determinant quantum Monte Carlo (DQMC) has been applied to finite-size lattice QED3 [XYXuU12018, wangDynamics2019, janssenConfinement2020], but its computational cost scales as $O(N_{\tau}V_{s}^{3})$, with $N_{\tau}$ the imaginary-time extent and $V_{s}$ the spatial volume. This cubic scaling, together with slow dynamics of local gauge updates, restricts simulations to small lattices and low precision. In critical phases such as the DSL, DQMC also suffers from severe autocorrelation effects [panSelf2025], further reducing efficiency. As a result, DQMC has not been able to access sufficiently large scales to settle questions about operator scaling dimensions or to confirm the conformal properties of QED3.
Hybrid quantum Monte Carlo (HQMC) offers an appealing alternative [Scalettar1987, Duane1987, Gupta1988, Beyl2018]. Unlike local-update DQMC, HQMC updates gauge field configurations globally using Hamiltonian dynamics, while the pseudofermion representation replaces explicit determinant evaluations by iterative solutions of linear systems. In principle this reduces the complexity from $O(N_{\tau}V_{s}^{3})$ to $O(N_{\tau}V_{s})$, and is inherently well suited to parallelization. Yet in earlier condensed matter applications, HQMC faced difficulties such as poor matrix conditioning and diverging iteration counts [Scalettar1987, White1988], which limited its practical advantages. More recently, HQMC has been successfully applied to spin-fermion models [patel2024strange, lunts2023non], where optimizations including preconditioned conjugate gradient solvers, matrix-free multiplications, and GPU acceleration enabled large-scale simulations. With these computational improvements, the non-Fermi-liquid scaling in the fermionic self-energy and bosonic propagators (space-time correlation functions) are observed. These encouraging developments motivate a dedicated HQMC approach to lattice QED3.
Inspired by this progress, in this paper we apply HQMC to simulate a $U(1)$ gauge field model coupled to $N_{f}=2$ fermions in (2+1)d, fully accelerated on GPUs. We scale the simulation size up to $N_{\tau}\times L\times L=660\times 66\times 66$, which represents a major improvement over previous DQMC studies of this model [XYXuU12018, wangDynamics2019, janssenConfinement2020] (typically limited to $200\times 20\times 20$), and exceeds by more than a factor of three the largest HQMC applications in spin-fermion models [patel2024strange, lunts2023non]. This advance is achieved through three problem-specific GPU optimizations:
  • a tailored preconditioner that stabilizes and accelerates the conjugate gradient solver,
@@ Introduction
  • customized CUDA kernels for matrix–vector multiplication exploiting the sparsity pattern of the fermion matrix, and
  • CUDA graph compilation to minimize kernel-launch overhead.
With these optimizations, the computational complexity is reduced from DQMC’s $O(N_{\tau}V_{s}^{3})$ scaling to nearly linear $O(N_{\tau}V_{s})$, allowing us to reach an unprecedented regime of system sizes. At this scale, we are able to investigate the correlation functions of various fermion bilinear operators and conserved currents with asymptotic accuracy. We find that spin and bond correlations share the same scaling dimension in the thermodynamic limit, consistent with their identification as adjoint bilinears of the emergent $SU(4)$ symmetry. The extracted scaling dimension $\Delta_{\rm adj}$ lies in the range 1.751.751.951.95, in agreement with the most recent analytical $\epsilon$-expansion results [di2017scaling]. We further observe that flux–flux correlations are governed by the conserved current dimension $\Delta_{J}=2$, leading to asymptotic long-time decay $\sim 1/\tau^{4}$. These results provide new, unbiased evidence for the conformal nature of the $U(1)$ DSL in QED3. Beyond these physical findings, our computational developments establish a scalable framework to explore DSL stability and its transitions to symmetry-breaking phases with significantly reduced computational cost.
I Results
@@
Figure 1: Lattice model of QED3. Fermions live on the sites of the lattice represented by blue dots, while gauge fields live on the bonds of the lattice represented by yellow dots. Fermion hops between nearest-neighbor sites with a phase $e^{i\phi_{ij}}$. The plot shows two adjacent square lattice layers, which are connected with temporal bonds represented as black dashed line. Gauge fields on these temporal bonds are set to be zero. Within a square lattice layer, the four bond colors represent the four families of bonds, which will be used in the checkerboard decomposition, as shown in Eq. (22); within each family, the fermion hopping terms represented by the bonds all mutually commute. In the bottom right cube, the circles on the facets denote the positive directions of the magnetic flux going out of the cube. The magnetic flux through a space-time plaquette $\Box$ is defined as $\theta^{\mu\nu}=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$, where $\mu,\nu\in\{x,y,\tau\}$ denote the plaquette orientation, ijij are along the positive direction, and $\phi_{ij}=-\phi_{ji}$.
Model
@@
The models we study are quantum electrodynamics (QED3) defined on a cubic space-time lattice (square spatial lattice), coupled with two-flavor fermions, as shown in Fig. 1 [XYXuU12018, wangDynamics2019]. Here, we start with the action of these lattice QED3 models: $S=S_{F}+S_{B}=\int_{0}^{\beta}d\tau\left(L_{F}+L_{B}\right)$, where $\beta$ is the inverse temperature and $\tau$ is the imaginary time. The first term
@@
In this paper, we focus on a dubbed non-compact non-compact $U(1)$ gauge field described by the Lagrangian